In the army kitchen there are four pots for every three pans for every 21 glasses. All together there are 476 of these items. How many pans are there?
a=pots; b=pans; c=glasses; So, 4a + 3b + 21c = 476
Now what?

Answer by Alan3354(69443)

In the army kitchen there are four pots for every three pans for every 21 glasses. All together there are 476 of these items. How many pans are there?
a=pots; b=pans; c=glasses; So, 4a + 3b + 21c = 476
Now what?
----------------
So, 4a + 3b + 21c = 476
The trick is, a, b & c are integers. You can't have 2/3 of a pot or pan.
-----------
4 + 3 + 21 = 28
476/28 = 17
-----
pots = 4*17 = 68
pans = 3*17 = 51
glasses = 21*17 = 357